OpenStudy (anonymous):

Solve the proportion 2x/2x+15=2/x

OpenStudy (lalaly):

2x*x = 2(2x+15) 2x^2 = 4x+30 2(x^2-2x-15) =0 2(x-5)(x+3)=0 x=5 or x=-3